OpenStudy (anonymous):

I am having trouble solving this: x^2+10x+25 / x^2 + 9x + 20

jhonyy9 (jhonyy9):

do you know how factoriz these ?

jhonyy9 (jhonyy9):

x^2 +10x +25 = x*x +5x+5x +5*5 = ? hope this will help you

jhonyy9 (jhonyy9):

the second try it you ok ?

jhonyy9 (jhonyy9):

courage

jhonyy9 (jhonyy9):

x^2 +9x +20 = x*x +4x +5x + 5*4 = ?

OpenStudy (anonymous):

Thanks It does help.

jhonyy9 (jhonyy9):

my pleasure good luck